OpenStudy (anonymous):

divide. assume that all variables in any denominator are nonzero. z^3/z^7

OpenStudy (hoblos):

\[\frac{ z^{a} }{ z^{b} } = z^{a-b}\]

OpenStudy (hoblos):

you may use also \[z^{-c} = \frac{ 1 }{ z^{c} }\]

OpenStudy (the_fizicx99):

\[z ^{-b} = \frac{ 1 }{ z^b }\]